# Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion

### v1

@@ 4.1 Datasets and Preprocessing
The MIMIC-III dataset Johnson et al. (2016) is a publicly available EHR dataset containing de-identified patients who were admitted to ICUs between 2001 and 2012. It includes medical discharge summaries, lab test results, chest x-ray images and more. Our analysis focuses on EHR data from approximately 27,000 patients including complete medical discharge summaries and lab test results. The MIMIC-IV dataset Johnson et al. (2023) comprises EHR data from 2008 to 2019. We utilize approximately 29,000 EHR records from MIMIC-IV, which include complete medical discharge summaries and lab test results. Our study targets 25 disease phenotypes as defined in the MIMIC-III benchmark Harutyunyan et al. (2019a).
@@
For medical notes, we extract the brief course from medical discharge summaries, removing numbers, noise, and stopwords. Numerical lab test results are converted into time series data using the benchmark tools Harutyunyan et al. (2019b), with missing values filled using the nearest available numbers. Time series anomaly descriptions are used the method defined in Appendix A.2. Data splitting follows the guidelines in Harutyunyan et al. (2019b), using a 4:1 ratio for training and testing.
@@ 4.2 Baseline Methods
We benchmark our approach against a range of methods: GRU Cho et al. (2014), PatchTST Nie et al. (2022), TimeLLM Jin et al. (2023), CAML Mullenbach et al. (2018), DIPOLE Ma et al. (2017), PROMPTEHR Wang and Sun (2022), LLaMA–7B Touvron et al. (2023), LDAM Niu et al. (2021a), FROZEN Tsimpoukelli et al. (2021), and EHR-KnowGen Niu et al. (2024). Detailed configurations of these baselines are provided in Appendix A.3. For the disease diagnosis task, we adopt two scales of Flan-T5 Chung et al. (2024) as the inference LLM to validate our model’s effectiveness in understanding multimodal EHRs, primarily driven by resource considerations and ease of experimentation. The Flan-T5-Small-based model is denoted as PromMedTS, while the Flan-T5-Large-based model is denoted as ProMedTS*. In principle, any sufficiently LLM could be substituted to potentially achieve even stronger results.
To ensure a fair comparison, all baselines also employ Flan-T5 as their backbone. Reported results are averaged over five runs with different random seeds. The statistical significance determined at p < 0.05 by t-test. Implementation details for every model are described in Appendix A.4, and training instructions appear in Appendix A.5. Our code is publicly available at https://anonymous.4open.science/r/ProMedTS-V1-5F51/.
@@ 4.3 Disease Diagnosis Performance
Table 1 shows that ProMedTS achieves the highest overall performance, particularly in F1 scores on MIMIC-IV. In addition, replacing the LLM with a larger model improves F1 scores on both datasets, indicating our model’s scalability and robustness across different LLMs. Furthermore, TimeLLM performs strongly with lab test, highlighting the value of time-series inputs for LLMs in disease diagnosis. Text-based methods (e.g., LLaMA) generally outperform time-series approaches, suggesting that medical notes capture richer disease-related information. Multimodal models (LDAM, FROZEN, EHR-KnowGen) exceed single-modality baselines (TimeLLM, CAML, DIPOLE, PROMPTEHR), confirming the benefits of integrating text and time series. Generative approaches (TimeLLM, LLaMA, FROZEN, EHR-KnowGen) also outperform classification-based methods. Although LLaMA performs well, its higher variance and parameter requirements reduce practicality. Notably, the Flan-T5-based ProMedTS and ProMedTS* surpass all baselines in F1 on both datasets, highlighting its efficiency and effectiveness.
@@ 4.4.1 Effect of Modality Alignment in ProMedTS
This section presents ablation studies to evaluate each module in ProMedTS. ProMedTS w/o LAB excludes lab test, removing modality alignment with anomaly descriptions and medical notes. ProMedTS w/o ANOMALY removes alignment with anomaly descriptions while keeping alignment between lab test and medical notes to assess the impact of self-supervision. Table 2 summarizes the results, showing that ProMedTS w/o LAB suffers a significant drop in F1 scores, highlighting the importance of lab test. ProMedTS w/o ANOMALY also shows reduced performance, highlighting the challenges of aligning modalities from discrete and continuous encoding spaces and the adverse effects of misalignment on multimodal understanding.
@@ 4.4.2 Impact of Self-Supervised Loss Functions
Table 3 summarizes an ablation study on the loss functions in ProMedTS. Both ProMedTS w/o CONTRAST and ProMedTS w/o MATCH show slight declines in F1 scores, emphasizing the importance of $\mathcal{L}_{contrast}$ for aligning and unifying time series and textual inputs within a shared latent space. The results also underscore the role of $\mathcal{L}_{match}$ in intra-modal alignment, ensuring the distinctiveness of time series data by aligning lab test with time series prompt embeddings. Notably, ProMedTS w/o GEN exhibits a significant drop in F1 scores, highlighting the critical role of $\mathcal{L}_{gen}$ in refining prompt embeddings and integrating temporal information from time series data and anomaly descriptions.
@@ 4.5 Model Efficiency and Complexity
Table 4: Sensitivity analysis on different length of time series prompt embedding.
4.6 Sensitivity Analysis of Time Series Prompt Length
We performed a sensitivity analysis to examine the impact of the time series prompt embedding length ($N_{p}$) on the performance of ProMedTS in disease diagnosis. Table 4 shows the F1 scores for embedding lengths of 12, 24, and 36. Slight fluctuations are observed in both micro and macro F1 scores across the MIMIC-III and MIMIC-IV datasets. The optimal embedding length is 24 for both datasets, consistent with the configuration used in our experiments.
@@ 4.6 Sensitivity Analysis of Time Series Prompt Length
Table 5: Micro and Macro F1 Scores Across Various Lab Test Input Types on Flan-T5 for Disease Diagnosis
4.7 Evaluating the Role of Anomaly Descriptions
@@ 5 Conclusion and Future Work
In this paper, we introduce ProMedTS, a lightweight and effective modality fusion framework that leverages self-supervised prompt learning for multimodal EHR integration. By bridging the modality gap between medical notes and lab test results, ProMedTS enables LLMs to process structured and unstructured medical data more effectively. Its three key modules and self-supervised loss functions advance language–time series integration in healthcare, providing a scalable and adaptable approach for real-world clinical applications. Evaluation on two EHR datasets demonstrates that ProMedTS significantly outperforms existing models in disease diagnosis, underscoring its potential to enhance clinical decision-making and improve patient care. In future work, we plan to extend our approach to larger and more diverse datasets, explore additional LLM architectures, and investigate further improvements in modality alignment techniques.
@@
While this study focuses on modality alignments and their application in downstream tasks, enhancing the explainability of disease diagnosis remains an area for future work, where we plan to incorporate the Chain-of-Thought rationale Wei et al. (2022). Additionally, computational constraints required the use of a relatively compact LLM, limiting the amount of clinical text processed at once, which may impact the model’s ability to leverage full medical histories. Expanding to more capable models will help address this challenge. Furthermore, our study primarily targets higher-level disease phenotypes in the International Classification of Diseases (ICD) codes Slee (1978), which could be expended to more downstream tasks. Future work will explore larger models, such as LLaMA Touvron et al. (2023) and Mistral Jiang et al. (2023), to improve diagnostic granularity and broaden coverage.
@@ Appendix A Appendix
A.1 Lab Test Anomaly Caption
Time series anomaly descriptions are generated using the IQR method Vinutha et al. (2018) to identify anomalies, capturing their timing and polarity (above or below standard values) and describing them with handcrafted templates. To caption the lab test anomaly into textual format, we design several text templates to describe the lab test anomalies. All templates are illustrated in Table 6.
A.2 Algorithm
@@
A.3 Baseline Models
If lab test value is not an abnormal value
@@ A.3 Baseline Models
If the lab test value is an abnormal value higher than the standard.
@@
If the lab test value is an abnormal value lower than the standard.
@@
If the lab test value is an abnormal that include both higher and lower than the standard value.
@@
    Diagnose disease from the following medical notes and lab test:
    Medical Notes: ms woman signficant pmh atrial fibrillation lung cancer resection congestive heart
    hypertension worsening diarrhea dysuria hypotensive admitted unit presumed active issues hypotensive
    pronounced leukocytosis multiple potential sources ct scan peritoneal fluids free fluid started surgery …
    Lab Test: $<prefix_{1}>,<prefix_{2}>,...,<prefix_{n}>$
    Diagnosis:
    Acute and unspecified renal failure, Fluid and electrolyte disorders, Septicemia (except in labor), Shock,
    Chronic obstructive pulmonary disease and bronchiectasis, Disorders of lipid metabolism,
    Cardiac dysrhythmias, Congestive heart failure; nonhypertensive, Diabetes mellitus with complications,
    Other liver diseases.
    Table 7: Training Instruction Template
@@
  • LDAM: LDAM Niu et al. (2021a) leverages multimodal inputs, combining laboratory testing results and medical notes for disease risk prediction. It utilizes label embedding to effectively integrate these two modalities.
@@
  • EHR-KnowGen: EHR-KnowGen Niu et al. (2024), touted as the state-of-the-art in EHR multimodal learning models, focuses on disease diagnosis generation. For this study, external domain knowledge is excluded to ensure a fair comparison.
Figure 5: Sensitivity analysis of varying ratios in loss function components.
@@ A.4 Implementation Details
In experiments, we utilized PyTorch framework version 2.0.1, operating on a CUDA 11.7 environment. We employed the AdamW optimizer with a starting learning rate of $1e^{-5}$ and a weight decay parameter of 0.05. Additionally, we implemented a warm-up strategy covering 10% of the training duration. Our experiments were conducted on high-performance NVIDIA Tesla V100 GPUs. Within the ProMedTS model, we used 24 time series prompt embeddings, each with a dimensionality of 768. The model’s hidden layer size was maintained at 768 for modality alignment and adjusted to 512 for downstream tasks. To standardize the time series data input, we padded all lab test results to a uniform length of 1000 time steps, allowing us to divide the data into 125 patches, with each patch containing 8 time steps. The Flan-T5 is fine-tuned on two MIMIC datasets Johnson et al. (2016, 2023) and then frozen for downstream tasks.

### v2

@@ 4.1 Datasets and Preprocessing
The MIMIC-III dataset Johnson et al. (2016) is a publicly available EHR dataset containing de-identified patients who were admitted to ICUs between 2001 and 2012. It includes medical discharge summaries, lab test results, chest x-ray images and more. Our analysis focuses on EHR data from approximately 27,000 patients, including complete medical discharge summaries and lab test results. The MIMIC-IV dataset Johnson et al. (2023) comprises EHR data from 2008 to 2019. We utilize approximately 29,000 EHR records from MIMIC-IV, which include complete medical discharge summaries and lab test results. Our study targets 25 disease phenotypes as defined in the MIMIC-III benchmark Harutyunyan et al. (2019a).
@@
For medical notes, we extract the brief course from discharge summaries, removing numbers, noise, and stopwords. Numerical lab tests are converted into time series data using the benchmark tools Harutyunyan et al. (2019b), with missing values filled using the nearest available numbers. Time series anomaly descriptions are used with the method defined in Appendix A.2. Data splitting follows the guidelines Harutyunyan et al. (2019b) using a 4:1 ratio for training and testing.
@@ 4.2 Baseline Methods
We benchmark our approach against a range of methods: GRU Cho et al. (2014), PatchTST Nie et al. (2022), TimeLLM Jin et al. (2023), CAML Mullenbach et al. (2018), DIPOLE Ma et al. (2017), Flan-T5 Chung et al. (2024), OPT Zhang et al. (2022), QWEN-2.5 Qwen et al. (2025), PROMPTEHR Wang and Sun (2022), LLaMA-1-7B Touvron et al. (2023) with anomalies input, LDAM Niu et al. (2021), FROZEN Tsimpoukelli et al. (2021), and EHR-KnowGen Niu et al. (2024a). Detailed configurations of these baselines are provided in Appendix A.3. For the disease diagnosis task, we adopt two scales of Flan-T5 Chung et al. (2024), OPT-0.1B, and QWEN-2.5-0.5B as the frozen LLM to validate our model’s effectiveness. To ensure a fair comparison, all baselines also employ Flan-T5-small as their backbone. Reported results are averaged over five runs with different random seeds. The statistical significance was determined at p < 0.05 by t-test. Implementation details for every model are described in Appendix A.4, and the example of training instructions appears in Appendix A.5. Our code is publicly available at github [^1] https://github.com/Healthcare-Data-Mining-Laboratory/PromptMedTS-V1.
@@ 4.3 Disease Diagnosis Performance
Table 1 shows the evaluation performance of disease diagnosis for our model ProMedTS with all baselines. First, for a single modality model, TimeLLM performs strongly with lab tests, highlighting the value of time-series inputs for LLMs in disease diagnosis. Text-based methods (e.g., Flan-T5) generally outperform time-series approaches, suggesting that medical notes capture richer disease-related information. Multimodal models (e.g., EHR-KnowGen, LLaMA) exceed single-modality baselines (e.g., TimeLLM, PROMPTEHR), confirming the benefits of integrating text and time series. Generative approaches (e.g., TimeLLM, LLaMA, EHR-KnowGen) also outperform classification-based methods. Although LLaMA performs well, its higher variance and parameter requirements reduce practicality.
Notably, the performance of our ProMedTS model, which employs Flan-T5-small as its backbone LLM, exceeds that of all baseline methods (especially LLaMA-1-7B) across both micro and macro F1 scores. This outcome emphasizes the efficiency and effectiveness of our approach in enabling LLMs to achieve multimodal understanding for disease diagnosis tasks. Additionally, ProMedTS consistently improves disease diagnosis performance across various backbone LLMs (Flan-T5-small&large, OPT, QWEN), achieving an average F1 score improvement of around 3%. These findings underscore our model’s scalability, robustness, and the generalizability of our approach.
@@ 4.4.1 Effect of Modality Alignment in ProMedTS
This section presents ablation studies to evaluate each module in ProMedTS. ProMedTS w/o LAB excludes lab test, removing modality alignment with anomaly descriptions and medical notes. ProMedTS w/o ANOMALY removes alignment with anomaly descriptions while keeping alignment between lab test and medical notes to assess the impact of self-supervision. Table 2 summarizes the results, showing that ProMedTS w/o LAB suffers a significant drop in F1 scores, highlighting the importance of the lab test. ProMedTS w/o ANOMALY also shows reduced performance, highlighting the challenges of aligning modalities from discrete and continuous encoding spaces and the adverse effects of misalignment on multimodal understanding. These findings emphasize the crucial role of both lab tests and anomaly captions as input for learning accurate time series embeddings, which in turn assist LLMs in disease diagnosis.
@@ 4.4.2 Impact of Self-Supervised Loss Functions
Table 3 summarizes an ablation study on the loss functions in ProMedTS. Both ProMedTS w/o CONTRAST and ProMedTS w/o MATCH show slight declines in F1 scores, emphasizing the importance of $\mathcal{L}_{contrast}$ for aligning and unifying time series and textual inputs within a shared latent space. The results also underscore the role of $\mathcal{L}_{match}$ in intra-modal alignment, ensuring the distinctiveness of time series data by aligning lab tests with time series prompt embeddings. Notably, ProMedTS w/o GEN exhibits a significant drop in F1 scores, highlighting the critical role of $\mathcal{L}_{gen}$ in refining prompt embeddings and integrating temporal information from time series data and anomaly descriptions.
@@ 4.5 Model Efficiency and Complexity
In addition, we conducted an efficiency evaluation using EHR-KnowGen with different backbone LLMs. As shown in Table 4, ProMedTS outperformed EHR-KnowGen by an average of 2% improvement in F1 scores while maintaining comparable training times with backbone LLMs of OPT and QWEN-2.5. This result supports the effectiveness and efficiency of our model in disease diagnosis.
4.6 Sensitivity Analysis of Time Series Prompt Length
We performed a sensitivity analysis to examine the impact of the time series prompt embedding length ($N_{p}$) on the performance of ProMedTS in disease diagnosis. Table 5 shows the F1 scores for embedding lengths of 12, 24, and 36. Slight fluctuations are observed in both micro and macro F1 scores across datasets. The optimal embedding length is 24 for both datasets, consistent with the configuration used in our experiments.
Models Training Micro F1 Macro F1
Time (h)
MIMIC-III
EHR-KnowGen
w/ OPT 8h 53.11 52.85
w/ QWEN-2.5 12h 54.09 51.45
ProMedTS
w/ OPT 8h 54.74 53.68
w/ QWEN-2.5 14h 57.82 55.97
MIMIC-III
EHR-KnowGen
w/ OPT 8.5h 52.46 50.19
w/ QWEN-2.5 13h 54.19 52.77
ProMedTS
w/ OPT 9h 55.10 53.23
w/ QWEN-2.5 14.5h 58.14 57.51
Table 4: Performance and Efficiency Comparison on Disease Diagnosis Tasks with Different LLMs.
@@ 4.6 Sensitivity Analysis of Time Series Prompt Length
Table 5: Sensitivity analysis on different lengths of time series prompt embedding.
4.7 Evaluating the Role of Anomaly Descriptions
@@
Table 6: Micro and Macro F1 Scores Across Various Lab Test Input Types on the LLM for Disease Diagnosis
@@ 5 Conclusion and Future Work
In this paper, we introduce ProMedTS, a lightweight and effective modality fusion framework that leverages self-supervised time series prompt learning for multimodal EHR integration. By bridging the modality gap between medical notes and lab test results, ProMedTS enables LLMs to process structured and unstructured medical data more effectively. Its three key modules and loss functions advance language–time series integration in healthcare, providing a scalable and adaptable approach for real-world clinical applications. Evaluation on two real-world EHR datasets demonstrates that ProMedTS significantly outperforms existing models in disease diagnosis, underscoring its potential to enhance LLMs’ clinical decision-making and improve patient care. In future work, we plan to extend our approach to apply to larger and more diverse datasets, such as Time-MMD Liu et al. (2024), conduct experiments with different clinical tasks, explore additional LLM architectures, and investigate further improvements in modality alignment techniques.
@@
While this study focuses on modality alignments and their application in downstream tasks, enhancing the explainability of disease diagnosis remains an area for future work, where we plan to incorporate the Chain-of-Thought rationale Wei et al. (2022) via knowledge distillation Hsieh et al. (2023); Lin et al. (2023). Furthermore, our study primarily targets higher-level disease phenotypes Slee (1978), which could be expanded to more downstream tasks.
@@
Acknowledgments
This work is supported by Tencent Rhino-Bird Focused Research Program (Value-aligned Credible Large Language Model) and RMGS project (Artificial Intelligence and Big Data Analytics for Social Good), and is also supported by the National Natural Science Foundation of China (No.62432006, No.62276159).
@@ Appendix A Appendix
A.1 Algorithm
@@
A.2 Lab Test Anomaly Caption
Time series anomaly descriptions are generated using the IQR method Vinutha et al. (2018) to identify anomalies, capturing their timing and polarity (above or below standard values) and describing them with handcrafted templates. To caption the lab test anomaly in textual format, we design several text templates to describe the lab test anomalies. All templates are illustrated in Table 7.
If lab test value is not an abnormal value:
@@ A.2 Lab Test Anomaly Caption
If the lab test value is an abnormal value higher than the standard:
@@
If the lab test value is an abnormal value lower than the standard:
@@
If the lab test value is an abnormal that include both higher and lower than the standard value:
@@
Diagnose disease from the following medical notes and lab test:
Medical Notes:
Diagnose disease from the following medical notes: woman decompensated etoh cirrhosis
initially fatigue ascites hydrothorax uti patient complicated hospital transferred icu hypoxemic
respiratory failure developed hypotension setting bleeding esophageal varices underwent tips
banding episodes gi bleeding … prevention outpatient electrolyte abnormalities refeeding syndrome…
contact narrow urine culture positive plan continue day add fosfomycin …
Lab test Anomaly Descriptions : $\leftarrow$ Only used during multimodal alignment
diastolic blood pressure is higher than normal one times, fraction inspired oxygen is higher than
normal forty-six times, glucose is higher than normal one times, heart rate is normal all the time,
mean blood pressure is normal all the time, oxygen saturation is normal all the time, respiratory rate
is higher than normal four times, systolic blood pressure is normal all the time, temperature is normal
all the time, weight is normal all the time, ph is normal all the time
Lab Test:
[[73. 1. 120. 87. 73. 94. 24. 105. 37. 81. 7.53 ],
[ 79. 1. 120. 87. 94. 94. 30. 131. 37. 81. 7.53 ],
[73. 1. 217. 87. 73. 94. 24. 105. 37. 81. 7.53 ],
[73. 1. 120. 87. 73. 94. 22. 105. 37. 81. 7.53 ]]
Diagnosis:
Diagnosed Results: Acute and unspecified renal failure, Fluid and electrolyte disorders,
Septicemia (except in labor), Shock, Chronic obstructive pulmonary disease and bronchiectasis,
Disorders of lipid metabolism, Cardiac dysrhythmias, Congestive heart failure; nonhypertensive,
Diabetes mellitus with complications, Other liver diseases.
Table 8: The Example of Training Instruction Template
Figure 5: Sensitivity analysis of varying ratios in loss function components, showing Micro and Macro F1 scores for downstream tasks.
A.3 Baseline Models
@@
  • Flan-T5: showcased within the scaling instruction-fine-tuning framework for language models Chung et al. (2024). It benefits from training on a wide array of datasets geared toward tasks like summarization and question answering.
  • OPT:OPT is a decoder-only model, designed to match the performance of GPT-3 Brown et al. (2020) while requiring only 1/7 of its computational resources. Trained primarily on around 180 billion tokens, OPT excels in tasks like poetry generation, dialogue, few-shot translation, and programming Zhang et al. (2022).
  • QWEN-2.5: QWEN-2.5 is a cutting-edge language model, expanding its pre-training dataset to 18 trillion tokens to enhance common sense, expert knowledge, and reasoning. It undergoes extensive supervised fine-tuning and multi-stage reinforcement learning, achieving top-tier performance in benchmarks for language understanding, reasoning, mathematics, coding, and human preference alignment Qwen et al. (2025).
@@
  • LDAM: LDAM Niu et al. (2021) leverages multimodal inputs, combining laboratory testing results and medical notes for disease risk prediction. It utilizes label embedding to effectively integrate these two modalities.
@@
  • EHR-KnowGen: EHR-KnowGen Niu et al. (2024a), touted as the state-of-the-art in EHR multimodal learning models, focuses on disease diagnosis generation. For this study, external domain knowledge is excluded to ensure a fair comparison.
@@ A.4 Implementation Details
In experiments, we utilized the PyTorch framework version 2.0.1, operating on a CUDA 11.7 environment. We employed the AdamW optimizer with a starting learning rate of $1e^{-5}$ and a weight decay parameter of 0.05. Additionally, we implemented a warm-up strategy covering 10% of the training duration. Our experiments were conducted on high-performance NVIDIA Tesla V100 GPUs. Within the ProMedTS model, we used 24 time series prompt embeddings, each with a dimensionality of 768. The model’s hidden layer size was maintained at 768 for modality alignment and adjusted to 512 for downstream tasks. To standardize the time series data input, we padded all lab test results to a uniform length of 1000 time steps, allowing us to divide the data into 125 patches, with each patch containing 8 time steps. All LLMs are fine-tuned on two MIMIC datasets Johnson et al. (2016, 2023) and then frozen for downstream tasks.